1.The required return on an investment is 8 percent. You estimate that firm A’s dividends will grow as follows:

Year

Dividend

1

$1.25

2

$1.89

3

$2.25

4

$3.00

For the subsequent years you expect the dividend to grow but at the modest rate of 3.5 percent annually. What is the maximum price that you should pay for this stock?

Please show work
Answer #1
Year Dividend
1 $                          1.2500
2 $                          1.8900
3 $                          2.2500
4 $                          3.0000
5 =$3*1.035 3.105
Terminal Value at the end of year 4 = Year 5 Dividend / (required rate- growth rate)
=$3.105/0.08-0.035
69.0000
Calculation of stock price
Year Cash Flow PV Factor PV Of Cash Flow
a b c=1/1.08^a d=b*c
1 $         1.2500 0.92593 $                     1.16
2 $         1.8900 0.85734 $                     1.62
3 $         2.2500 0.79383 $                     1.79
4 $         3.0000 0.73503 $                     2.21
4 $            69.00 0.73503 $                   50.72
Stock Price $                   57.49